OpenStudy (anonymous):

A line has a slope of 2 and passes through the point (4, 0). What are the coordinates of another point that is on the line? a.(2, 0) b.(0, 2) c.(4, 6) d.(6, 4)

jimthompson5910 (jim_thompson5910):

First you have to determine the equation of the line that has a slope of 2 and goes through the point (4, 0).

jimthompson5910 (jim_thompson5910):

Do you know how to do this?

OpenStudy (anonymous):

No >.<

jimthompson5910 (jim_thompson5910):

The general linear equation is y = mx+b where m is the slope so m = 2 which means y = mx+b becomes y = 2x+b

jimthompson5910 (jim_thompson5910):

We're given the point (4,0), so x = 4 and y = 0

jimthompson5910 (jim_thompson5910):

Plug these two in to get y = 2x+b 0 = 2(4)+b 0 = 8 + b Keep going to solve for b